# ARMAr-LASSO: Mitigating the Impact of Predictor Serial Correlation on the LASSO

### v2

@@ Appendix F Analysis of the minimum eigenvalues Under DGPs(A)-(D)
In this section, we compare the minimum eigenvalues of the design matrix of LAS, GLS-LAS, and ARMAr-LAS in the case of $n=50$. Figure S.10 shows the average of the minimum eigenvalues obtained in the experiments presented in Section 4.1.1. Both LAS and GLS-LAS reduce their minimum eigenvalues as $\phi$ increases. This does not happen for ARMAr-LAS, which maintains the same value regardless of the degree of serial correlation. Figure S.11 shows the same results but for the experiments presented in Section 4.1.2. In this case, we compare the minimum eigenvalues for the two DGPs (reported as 0 for DGP (C) and 1 for DGP (D)). Again, ARMAr-LAS maintains larger minimum eigenvalue with respect to LAS and GLS-LAS.
@@
Figure S.11: Minimum eigenvalues for the design matrix of LAS, GLS-LAS, and ARMAr-LAS, for DGPs (C) (0) and (D) (1).
Appendix G Performance in a Large $T$ Regime
Here we compare our ARMAr-LAS with the employed LASSO-based benchmarks in the case of DGP (A) (see Section 4.1.1) with $T=1500$, $n=50$, and SNR=10. This section aims to evaluate the performances of ARMAr-LAS in a large sample size regime. Results in Table S.6 show that ARMAr-LAS performs as GLS-LAS. This result is expected since under DGP (A) these two estimators coincide. Further, both outperform the other LASSO-based methods providing more accurate coefficient estimates and forecasts, as well as a perfect variable selection accuracy.
@@ Appendix G Performance in a Large $T$ Regime
Appendix H Performance with Misspecified Autoregressive Structure
In this section, we compare our ARMAr-LASSO with the LASSO-based benchmarks in the case where the former misspecifies the autoregressive model of predictors. In particular, we generated both predictors and error terms from an AR(2) model with autoregressive coefficients equal to 1.2 and -0.4, but predictors are filtered through an AR(1) model by ARMAr-LAS. We consider $T=150$, SNR=10, and $n=50,150,300$. Results are reported in Table S.7. Also in this case where the predictors are filtered with a misspecified autoregressive model, ARMAr-LAS outperforms LASSO-based benchmarks. This is because, despite the misspecification, the filter can remove the majority of serial correlation with the proper estimation of a single autoregressive coefficient. This is corroborated by the averages of the minimum eigenvalues of the correlation matrices for LAS, GLS-LAS, and ARMAr-LAS which are 0.00680, 0.01824, and 0.02861, respectively.
@@ Appendix H Performance with Misspecified Autoregressive Structure
Appendix I List of Time Series in the Euro Area Data
We report the list of series for the Euro Area dataset adopted in the forecasting exercise (obtained from Proietti and Giovannelli (2021)). As for the FRED data, the column tcode denotes the data transformation for a given series $x_{t}$: (1) no transformation; (2) $\Delta x_{t}$; (3)$\Delta^{2}x_{t}$; (4) $\text{log}(x_{t})$; (5) $\Delta\text{log}(x_{t})$; (6) $\Delta^{2}\text{log}(x_{t})$. (7) $\Delta(x_{t}/x_{t-}-1.0)$.
@@ Appendix I List of Time Series in the Euro Area Data
Table S.8: Euro Area macroeconomic variables from Proietti and Giovannelli (2021)
@@ 84 Svy - Households, Economic Situation Next 12M FR CCI 1
[…]
@@
Appendix J Distribution of $b$
Consider two orthogonal Gaussian AR(1) processes generated according to the model ${x}_{i,t}=\phi_{i}{x}_{i,t-1}+{u}_{i,t}$, where $u_{i,t}\sim N(0,1)$, $i=1,2,\ t=1,\dots,100$ and $\phi_{1}=\phi_{2}=\phi$. In this simulation exercise we run the model
$x_{2t}=\beta x_{1t}+e_{t},$ (J.14)
where $e_{t}\sim N(0,\sigma_{e}^{2})$, and study the distribution of the OLS estimator $b$ of $\beta$ in the following four cases in terms of degrees of serial correlation: $\phi=0.3,0.6,0.9,0.95$. Figure S.12 reports the density of $b$ across the $\phi$ values obtained on 5000 Monte Carlo replications. We compare this density with that of three zero-mean Gaussian variables where the variances are respectively:
  • (a)
    $S_{1}^{2}=\frac{\widehat{\sigma}^{2}_{\widehat{e}}}{\sum_{t=1}^{T}(x_{1t}-\overline{x}_{1})^{2}}$, where $\widehat{\sigma}^{2}_{\widehat{e}}$ is the sample variance of the estimated residual $\widehat{e}_{t}=x_{2t}-bx_{1t}$. This is the OLS estimator for the variance of $\beta$.
  • (b)
    $S_{2}^{2}=\frac{1}{T}\frac{\frac{1}{T-2}\sum_{t=1}^{T}(x_{1t}-\overline{x}_{1})^{2}\widehat{e}_{t}^{2}}{\mathopen{}\mathclose{{}\left[\frac{1}{T}\sum_{t=1}^{T}(x_{1t}-\overline{x}_{1})^{2}}\right]^{2}}\widehat{f}_{t}$, is the Newey-West (NW) HAC estimator (Newey and West, 1987), where $\widehat{f}_{t}=\mathopen{}\mathclose{{}\left(1+2\sum_{j=1}^{m-1}\mathopen{}\mathclose{{}\left(\frac{m-j}{m}}\right)\widehat{\rho}_{j}}\right)$ is the correction factor that adjusts for serially correlated errors and involves estimates of $m-1$ autocorrelation coefficients $\widehat{\rho}_{j}$, and $\widehat{\rho}_{j}=\frac{\sum_{t=j+1}^{T}\widehat{v}_{t}\widehat{v}_{t-j}}{\sum_{t=1}^{T}\widehat{v}_{t}^{2}}$, with $\widehat{v}_{t}=(x_{1t}-\overline{x}_{1})\widehat{e}_{t}$. A rule of thumb for choosing $m$ is $m=[0.75T^{1/3}]$.
  • (c)
    $S_{3}^{2}=\frac{(1-\phi_{1}^{2}\phi_{2}^{2})(1-\phi_{1}^{2})}{(T-1)(1-\phi_{2}^{2})(1-\phi_{1}\phi_{2})^{2}}$, is the theoretical variance of $b$ obtained in Lemma B.1.
From Figure S.12 we observe that the variance of $b$ increases with the degree of serial correlation ($\phi$) in a way that is well approximated by the distribution derived in Lemma B.1 (see dotted line). On the contrary, OLS (solid line) and NW (dashed line), are highly sub-optimal in the presence of strong serial correlation, underestimating the variability of $b$ as the serial correlation increases.
(a) $\phi=0.3,\ T=100$
(b) $\phi=0.6,\ T=100$
(c) $\phi=0.9,\ T=100$
(d) $\phi=0.95,\ T=100$
Figure S.12: Density of $b$ between uncorrelated AR(1) Gaussian processes. Solid line indicates the approximated density obtained by using the classical OLS estimator, dashed line indicates the approximated density obtained by using the NW estimator, and, finally, dotted line shows the theoretical approximated density obtained in Lemma B.1.
Appendix K Distribution of $\widehat{Cov}(u_{1},u_{2})$
In Figure S.13 we report the density of $\widehat{Cov}(u_{1},u_{2})$ when $u_{1}$ and $u_{2}$ are standard Normal in the cases of $T=30$ and 100. Red line shows the density of $N\mathopen{}\mathclose{{}\left(0,\frac{1}{T-1}}\right)$. Observations are obtained on 5000 Monte Carlo replications. We observe that the approximation of $\widehat{Cov}(u_{1},u_{2})$ to $N(0,\frac{1}{T-1})$ holds also when $T$ is small (see Figure S.13 (a) relative to $T$=30). In particular, for $T=30$, the p-value of the Shapiro test is 0.89, the skewness is 0.031 and the kurtosis is 3.001. For $T=100$, the values for the same statistics are 0.200, -0.016, and 3.146, respectively. This analysis corroborate numerically the results in Glen et al. (2004), which show that if $x$ and $y$ are $N\mathopen{}\mathclose{{}\left(0,1}\right)$, then the probability density function of $xy$ is $\frac{K_{0}(|xy|)}{pi}$, where $K_{0}(|xy|)$ is the Bessel function of the second kind.
(a) $T=30$
(b) $T=100$

### v3

@@ D.3 Analysis of the minimum eigenvalues
In this section, we compare the minimum eigenvalues of the design matrix of LAS, GLS-LAS, and ARMAr-LAS in the case of $n=50$ and $SNR=10$. Figure S.13 shows the average of the minimum eigenvalues obtained in the experiments presented in Section D.2. Both LAS and GLS-LAS reduce their minimum eigenvalues as $\phi$ increases. This does not happen for ARMAr-LAS, which maintains the same value regardless of the degree of serial correlation. Figure S.14 shows the same results but for the experiments presented in Section 4.1.2 in the main text. In this case, we compare the minimum eigenvalues for the two DGPs reported as 0 for no common factor ($q=0$) and 1 for common factor ($q=1$). Again, ARMAr-LAS maintains larger minimum eigenvalue with respect to LAS and GLS-LAS. This analysis corroborates the statement of Remark 6 of the main text.
@@
Figure S.14: Minimum eigenvalues for the design matrix of LAS, GLS-LAS, and ARMAr-LAS, for DGP in Section 4.1.2 of the main text. In this case, we compare the minimum eigenvalues for the two DGPs reported as 0 for no common factor ($q=0$) and 1 for common factor ($q=1$).
D.4 Performance in a Large $T$ Regime
Here we compare our ARMAr-LAS with the employed LASSO-based benchmarks in the case of DGP (A) (see Section D.2) with $T=1500$, $n=50$, and SNR=10. This section aims to evaluate the performances of ARMAr-LAS in a large sample size regime. Results in Table S.5 show that ARMAr-LAS performs as GLS-LAS. This result is expected since under DGP (A) these two estimators coincide. Further, both outperform the other LASSO-based methods providing more accurate coefficient estimates and forecasts, as well as a perfect variable selection accuracy.
@@ D.4 Performance in a Large $T$ Regime
D.5 Performance with Misspecified Autoregressive Structure
In this section, we compare our ARMAr-LASSO with the LASSO-based benchmarks in the case where the former misspecifies the autoregressive model of predictors. In particular, we generated both predictors and error terms from an AR(2) model with autoregressive coefficients equal to 1.2 and -0.4, but ARMAr-LASSO filters predictors through an AR(1) model. We consider $T=150$, SNR=10, and $n=50,150,300$. Results are reported in Table S.6. Also in this case where the predictors are filtered with a misspecified autoregressive model, ARMAr-LAS outperforms LASSO-based benchmarks. This is because, despite the misspecification, the filter can remove the majority of serial correlation with the proper estimation of a single autoregressive coefficient. This is corroborated by the averages of the minimum eigenvalues of the correlation matrices for LAS, GLS-LAS, and ARMAr-LAS which are 0.00680, 0.01824, and 0.02861, respectively.
@@ D.5 Performance with Misspecified Autoregressive Structure
Appendix E List of Time Series in the Euro Area Data
[…]
@@ Appendix E List of Time Series in the Euro Area Data
Table S.7: Euro Area macroeconomic variables from Ale2021
@@ 84 Svy - Households, Economic Situation Next 12M FR CCI 1
185 Euro Interbank Offered Rate - 3-Month (Mean) ($\boldsymbol{acf:\ 0.67}$) EA M&IR 5
@@
Appendix F Five most frequently selected predictors
Table S.8 reports the list of the top 5 predictors in terms of selection frequency across forecasting samples obtained from the empirical application in Section 4.2 of the main text. Regardless of the forecasting horizon $h$, the top predictor for ARMAr-LAS is the Goods Index. The other top predictors, also in the HCPI domain, include EA measurements (e.g., Services Index), or are specific to France and Germany (e.g., All-Items).
Table S.8: Five most frequently selected predictors. Selection percentages are ratios between the times a predictor appears in a forecast and the total number of forecasts (120 for hh=12 and 96 for $h=24$).
Rank Selected Variables
hh=12 $h=24$
Goods, Index Goods, Index
85.8% 85.4%
II° Industrial Goods, Index Services, Index
47.5% 43.8%
III° Services, Index All-Items (De)
40.8% 35.4%
32.5% 32.3%
All-Items (Fr) Industrial Goods, Index
24.2% 30.2%
Appendix G Distribution of $\widehat{Cov}(u_{1},u_{2})$
In Figure S.15 we report the density of $\widehat{Cov}(u_{1},u_{2})$ when $u_{1}$ and $u_{2}$ are standard Normal in the cases of $T=30$ and 100. Red line shows the density of $N\mathopen{}\mathclose{{\left(0,\frac{1}{T-1}}}\right)$. Observations are obtained on 5000 Monte Carlo replications. We observe that the approximation of $\widehat{Cov}(u_{1},u_{2})$ to $N(0,\frac{1}{T-1})$ holds also when $T$ is small (see Figure S.15 (a) relative to TT=30). In particular, for $T=30$, the p-value of the Shapiro test is 0.89, the skewness is 0.031 and the kurtosis is 3.001. For $T=100$, the values for the same statistics are 0.200, -0.016, and 3.146, respectively. This analysis corroborate numerically the results in Glen2004, which show that if $x$ and $y$ are $N\mathopen{}\mathclose{{\left(0,1}}\right)$, then the probability density function of $xy$ is $\frac{K_{0}(|xy|)}{pi}$, where $K_{0}(|xy|)$ is the Bessel function of the second kind.